4. A project to develop a county park has an actual cost in month 17 of $350,000, a planned cost of $475,000, and a value completed of S300,000. Find the cost and schedule variances and the three indexes.
Answer #1

Answer:

Actual cost or AC = $350,000,
Planned cost or PV = $475,000,
Value completed or earned value (EV) = $300,000
Cost Variance (CV) = 300,000 – 350,000
                                  = -50,000
Schedule Variance(SV) = 300,000 – 475,000
                                       = -175,000
Cost Performance Index (CPI) = Earned value/Actual cost
                                       = 300,000/350,000
                                        = 0.8571
Schedule Performance Index (SPI) = Earned value/Planned Cost
                                                  = 300,000/475,000
                                                  = 0.6316
Scheduled Time Work Performed (STWP)
Time ( t) = Schedule Variance/Slope of Planned costs = -175,000/ (475,000/17)
                                                      = - 6.2631months   
                            Time Difference      = 17- 6.2631   = 10.7369
                            Time Variance = 10.7369/17 = 0.6316